# SBI-RAG: Enhancing Math Word Problem Solving for Students through Schema-Based Instruction and Retrieval-Augmented Generation

### v1

@@ Abstract
Many students struggle with math word problems (MWPs), often finding it difficult to identify key information and select the appropriate mathematical operations. Schema-based instruction (SBI) is an evidence-based strategy that helps students categorize problems based on their structure, improving problem-solving accuracy. Building on this, we propose a Schema-Based Instruction Retrieval-Augmented Generation (SBI-RAG) framework that incorporates a large language model (LLM). Our approach emphasizes step-by-step reasoning by leveraging schemas to guide solution generation. We evaluate its performance on the GSM8K dataset, comparing it with GPT-4 and GPT-3.5 Turbo, and introduce a "reasoning score" metric to assess solution quality. Our findings suggest that SBI-RAG enhances reasoning clarity and problem-solving accuracy, potentially providing educational benefits for students.
@@ 3 Evaluation
Schema Classifier Results: The schema classifier was trained to identify two schema categories and three sub-categories: Additive Change, Additive Difference, Additive Total, Multiplicative Comparison, Multiplicative Equal Groups, and Multiplicative Ratios/Proportions. It achieved high precision, recall, and F1 scores, with an overall accuracy of 97%. The training and validation losses show consistent convergence, indicating effective learning without overfitting, ensuring reliable schema predictions across various problem types.
@@
LLM-as-a-Judge Results: We implemented the LLM-as-a-Judge approach [36] to evaluate the quality of reasoning in the responses generated by both our Schema-Based RAG system and the baseline LLMs. This method allows for an objective, scalable evaluation by approximating human judgment through the use of LLMs. Our LLM-as-a-Judge process involves scoring responses based on clarity, logical progression, and completeness. Results showed that the Schema-Based RAG approach consistently outperformed GPT-4 and GPT-3.5 Turbo in terms of reasoning quality. More details about the implementation and results can be found in Appendix G.
@@ D.4 Evaluation and Results
Figure 4: Confusion matrix for the schema classifier
Figure 5: Training and validation losses for the schema classifier

### v2

@@ Abstract
Many students struggle with math word problems (MWPs), often finding it difficult to identify key information and select the appropriate mathematical operations. Schema-based instruction (SBI) is an evidence-based strategy that helps students categorize problems based on their structure, improving problem-solving accuracy. Building on this, we propose a Schema-Based Instruction Retrieval-Augmented Generation (SBI-RAG) framework that incorporates a large language model (LLM). Our approach emphasizes step-by-step reasoning by leveraging schemas to guide solution generation. We evaluate its performance on the GSM8K dataset, comparing it with GPT-4 and GPT-3.5 Turbo, and introduce a "reasoning score" metric to assess solution quality. Our findings suggest that SBI-RAG enhances reasoning clarity and facilitates a more structured problem-solving process potentially providing educational benefits for students.
@@ 3 Evaluation
Schema Classifier Results: The schema classifier was trained to identify two schema categories and three sub-categories: Additive Change, Additive Difference, Additive Total, Multiplicative Comparison, Multiplicative Equal Groups, and Multiplicative Ratios/Proportions. As seen in Figure 3 and Figure 3, it achieved high precision, recall, and F1 scores, with an overall accuracy of 97%. The training and validation losses show consistent convergence, indicating effective learning without overfitting, ensuring reliable schema predictions across various problem types.
Figure 2: Confusion matrix for the schema classifier
Figure 3: Training and validation losses for the schema classifier
@@
LLM-as-a-Judge Results: We implemented the LLM-as-a-Judge approach [36] to evaluate the quality of reasoning in the responses generated by both our Schema-Based RAG system and the baseline LLMs. This method allows for an objective, scalable evaluation by approximating human judgment through the use of LLMs. Our LLM-as-a-Judge process involves scoring responses based on clarity, logical progression, and completeness. Results showed that the Schema-Based RAG approach consistently outperformed GPT-4 and GPT-3.5 Turbo in terms of reasoning quality.For more details refer to Appendix G.
@@ Appendix D Training Details and Implementation
The code for this implementation can be found on GitHub: https://github.com/pdx97/SBI-RAG_Neurips2024.
@@ Appendix G LLM-as-a-Judge Results and Task Definition
Additionally, For a given Question (As seen in Figure9), our evaluation using the LLM-as-a-Judge approach assessed responses based on three key sub-metrics: Clarity, Logical Progression, and Completeness as shown in Figure 10. The Total rating was then calculated based on these sub-metrics, which are defined as follows:
Figure 9: Sample question with Response 1 from SBI-RAG and Response 2 from GPT-4
  • Clarity: Assesses how clearly the response conveys the solution. A high clarity score indicates ease of understanding, appropriate language use, and avoidance of unnecessary jargon or complexity that could confuse the reader.
  • Logical Progression: Evaluates the logical flow of the response. A high score here indicates that each step follows naturally from the previous one, forming a coherent sequence that effectively guides the reader through the problem-solving process.
  • Completeness: Measures whether the response fully addresses all aspects of the question. A complete response includes all necessary steps, explanations, and justifications required to reach the solution.
Response Clarity (0-10) Logical Progression (0-10) Completeness (0-10) Total Rating (0-10)
Response 1 9.0 9.0 9.0 9.0
Response 2 8.0 7.5 7.0 7.5
Table 2: Evaluation Scores for Response 1 and Response 2
Figure 10: Overall Scores by LLM-as-a-Judge